# Uncertainty Quantification in Retrieval Augmented Question Answering

### v1

@@ B.3 Prompts
The prompt we use for our QA models is shown in Table 6. Table 7 illustrates the prompt used for our the p(true) baseline. Table 8 shows the prompt used for the LLM-based accuracy (AccLM) metric.
Appendix C Additional Results
C.1 Generalisation of Uncertainty Estimation
In this series of experiments, we assess the generalisation ability of our Passage Utility estimator. To this end, following previous work on question answering and out-of-distribution (o.o.d) scenarios (Kamath et al., 2020; Zhang et al., 2021); we train a Passage Utility predictor on the SQuAD dataset and then use it to predict zero-shot (i.e., without further fine-tuning) passage utilities on all other datasets’ test cases. As p(true) relies on 20 in context training examples, we also evaluate its ability to generalise to out of distribution test cases.
Table 9 shows AUROC for answer uncertainty prediction on o.o.d scenarios. We also report PPL as a baseline and the AUROC values for the in distribution scenario as upper-bound. Although Passage Utility performance decreases in o.o.d settings, it remains competitive in three out of five datasets. On NQ and WebQ the performance is slightly above the PPL baseline. Note that we focus on zero-shot accuracy to assess bare transfer performance; however, it would make sense to adapt the model with few examples from the o.o.d data (Kamath et al., 2020; Zhang et al., 2021). Interestingly, p(true)’s performance also drops in all o.o.d test sets showing that relying on a fixed number of training examples is neither robust nor has a principled and scalable adaptation method (e.g., fine-tuning).
NQ TQA WebQ PopQA RefuNQ
PPL 0.64 0.68 0.52 0.59 0.51
p(true) (i.i.d) 0.73 0.75 0.67 0.81
Passage Utility (i.i.d) 0.72 0.82 0.70 0.84
p(true) (o.o.d) 0.67 0.63 0.63 0.72 0.62
Passage Utility (o.o.d) 0.66 0.82 0.57 0.73 0.81
Table 9: Out-of-domain performance of Passage Utility predictor (with Gemma2-9B). Uncertainty predictors are trained on SQuAD and evaluated zero-shot on NQ, TQA, WebQ, PopQA, and RefuNQ test sets.
C.2 Reference Retrieval Augmented QA Accuracy
Table 10 shows retrieval augmented QA performance (Acc and AccLM) for the five QA models on the development and test sets of the six QA tasks.
C.3 Detailed Uncertainty Estimation Results
@@
Table 11 shows performance of uncertainty quantification approaches on the development sets. We report AUROC and AURAC.
@@
Acc AccLM Acc AccLM Acc AccLM Acc AccLM Acc AccLM Acc AccLM
Development
Gemma2.9B 0.48 0.66 0.74 0.80 0.46 0.66 0.38 0.60 0.51 0.52
@@
Gemma2.2B 0.43 0.59 0.67 0.73 0.47 0.65 0.34 0.55 0.48 0.49
Gemma2.27B 0.48 0.66 0.75 0.81 0.49 0.67 0.38 0.60 0.52 0.52
Test
Gemma2.9B 0.49 0.65 0.74 0.80 0.40 0.66 0.43 0.60 0.50 0.52 0.26 0.40
@@
Gemma2.2B 0.44 0.57 0.67 0.72 0.39 0.61 0.39 0.56 0.48 0.49 0.24 0.33
Gemma2.27B 0.48 0.65 0.76 0.81 0.41 0.66 0.42 0.61 0.51 0.53 0.26 0.39
Table 10: Performance of target QA models (with $|R|=5$) on the development and test sets. We report token- and model-based accuracy (Acc and AccLM). AccLM is computed by Qwen2-72B-Instruct.
NQ TQA WebQ SQuAD PopQA NQ TQA WebQ SQuAD PopQA
@@
Gemma2-9B
@@
Avg.Ans.Len 0.61 0.64 0.65 0.63 0.68 0.68 0.83 0.71 0.65 0.61
Passage Utility 0.71 0.83 0.71 0.83 0.86 0.74 0.88 0.75 0.76 0.72
Llama-3.1-8B
@@
Semantic Entropy 0.74 0.83 0.69 0.74 0.81 0.75 0.86 0.71 0.71 0.68
Avg.Ans.Len 0.73 0.73 0.69 0.69 0.84 0.73 0.82 0.71 0.67 0.69
Passage Utility 0.79 0.84 0.76 0.82 0.85 0.77 0.87 0.76 0.74 0.70
Mistral-7B-v0.3
@@
Avg.Ans.Len 0.66 0.75 0.65 0.68 0.81 0.69 0.85 0.73 0.67 0.70
Passage Utility 0.77 0.81 0.74 0.83 0.84 0.74 0.87 0.79 0.74 0.71
Table 11: Answer uncertainty estimation for QA models Gemma2-9B, Llama-3.1-8B, and Mistral-7B-v0.3 on NQ, TQA, WebQ, SQuAD, and PopQA development sets. We report AUROC and AURAC.
Appendix D Examples of False Positives and Negatives
Tables 1215 illustrate the working of Passage Utility for answer uncertainty estimation. As we report AUROC scores, we do not set any correct/incorrect decision threshold; for the purpose of this discussion, we assume a decision point at 0.5 and analyse clear success and failure cases. For each example, we show the question, gold, and generated answers in the top block. Then, we show three retrieved passages with their estimated Passage Utility and a final block with the ten sampled answers, their grouping into clusters, and the Cluster Assignment entropy.
Table 12 shows an example for a SQuAD question and the Llama-3.1-8B QA model. In this case, the QA model correctly answers and the Passage Utility estimate is high (i.e., indicating correct answer). Table 13 illustrates a case where Llama-3.1-8B’s answer is incorrect and all Passage Utilities are very low (i.e., indicating incorrect answer). The example from NQ in Table 14 shows a case where all Passage Utilities are low but the QA model (Gemma2-9B) answers correctly. The first passage is not useful, the second does not explicitly mention the answer but still primes the QA model to answer correctly, while the third passage mentions the answer.
@@ Appendix D Examples of False Positives and Negatives
Question: Due to increased unemployment, who mainly opposed the Shah’s regime? Generated Answer: Millions of youth and poor migrants. Gold Answer: Millions of youth who had migrated to the cities.
Youth who migrated to cities for construction jobs. Iran. unemployment, especially among millions of youth who had migrated to the cities of Iran looking for construction jobs during the boom years of the early 1970s. By the late 1970s, many of these people opposed the Shah’s regime and began to organize and join the protests against it. The 1979 Revolution, later known as the ‘‘Islamic Revolution’’, began in January 1978 with the first major demonstrations against the Shah. [cont.] 3.93
Unemployed and the poor. Ruhollah Khomeini. unemployment, ideological disagreement over the economy, and ‘‘international pressure and isolation’’ such as US sanctions following the hostage crisis. Due to the Iran–Iraq War, poverty is said to have risen by nearly 45% during the first 6 years of Khomeini’s rule. Emigration from Iran also developed, reportedly for the first time in the country’s history. Since the revolution and war with Iraq, an estimated ‘‘two to four million entrepreneurs, professionals, technicians, and skilled craftspeople (and their capital)’’ have emigrated to other countries. [cont.] -1.37
The National Front. Mohammad Reza Pahlavi. professors issued a public statement criticising the 1953 coup, all were dismissed from their jobs, but in the first of his many acts of ‘‘magnanimity’’ towards the National Front, Mohammad Reza intervened to have them reinstated. Mohammad Reza tried very hard to co-opt the supporters of the National Front by adopting some of their rhetoric and addressing their concerns, for example declaring in several speeches his concerns about the Third World economic conditions and poverty which prevailed in Iran, a matter that had not much interested him before. [cont.] -3.48
[’Migrants to the cities.’], [’Millions of youth who had migrated to cities.’, ’Millions of youth who migrated to cities.’], [’Cultural and religious conservatives, and the urban poor who had migrated to cities for jobs.’], [’Youth who had migrated to the cities.’], [’Millions of young migrants who had moved to cities in the early 1970s.’], [’Millions of youth and poor migrants to cities.’], [’Cultural and religious conservatives, mostly urban migrants.’, ’Cultural and religious conservatives, particularly the migrants to cities.’, ’Cultural and religious conservatives, especially those recently migrated to the cities.’] 1.83
@@
Question: Which company was targeted by the NAACP for not having fair practices? Generated Answer: Target Corporation. Gold Answer: Lockheed Aircraft Corporation.
Target Corporation. Target Corporation. of Colored People has repeatedly given Target failing grades on its annual Economic Reciprocity Initiative report card, a measure of the company’s ‘‘commitment to the African-American citizenry’’. In 2003 and 2005, the NAACP has rated Target an ‘‘F’’ on this report; in 2004, Target was rated a ‘‘D-’’. In 2006, when Target was asked why it didn’t participate in the survey again, a representative explained, ‘‘Target views diversity as being inclusive of all people from all different backgrounds, not just one group.’’ In February 2006, the National Federation of the Blind (NFB) filed a class action [cont.] -3.59
None, the NAACP was involved in the Duke lacrosse case. Reactions to the Duke lacrosse case. formed an opinion on the case. North Carolina NAACP Legal Redress Chair, Al McSurely, explained that ‘‘The NAACP stands for fair play for all parties, zealous investigation and deep concern for the survivors of racist/sexist attacks.’’ At the same time, some have criticized the NAACP for making statements that portrayed the players as racist despite evidence to the contrary, using the case to promote the group’s cause, and implying guilt. McSurely stated that ‘‘[w]ithin five minutes, the men threatened the women with racial and misogynist verbal assaults, [cont.] -3.67
Philadelphia Transit Company (PTC). Philadelphia transit strike of 1944. (PRTEU), Frank Carney, proved to be equally reticent and claimed that he was not authorized by the union members to consider a request to allow promotions of black employees. The black PTC employees enlisted the help of the NAACP and started lobbying the federal authorities, particularly the Fair Employment Practices Commission (FEPC), to intervene. The Fair Employment Practices Commission, created by an executive order of the President in 1941, was charged with ensuring non-discrimination employment practices by government contractors. [cont.] -3.49
@@
Table 13: True positive example from (SQuAD development set): Passage Utility predicts the answer provided by the QA model (Llama-3.1-8B) which is incorrect.
Question: Close Encounters of the Third Kind Oscar nominations? Generated Answer: Eight Oscar nominations. Gold Answer: Eight Oscars.
Sound effects editing. Close Encounters of the Third Kind. although the Academy honored the film’s sound effects editing with a Special Achievement Award (Frank Warner). At the 32nd British Academy Film Awards, ‘‘Close Encounters’’ won Best Production Design, and was nominated for Best Film, Direction, Screenplay, Actor in a Supporting Role (François Truffaut), Music, Cinematography, Editing, and Sound. ‘‘Close Encounters’’ lost the Hugo Award for Best Dramatic Presentation to ‘‘Star Wars’’ [cont.] -1.86
Eight Oscar nominations. Close Encounters of the Third Kind. Close Encounters of the Third Kind Close Encounters of the Third Kind is a 1977 American science fiction film written and directed by Steven Spielberg, and starring Richard Dreyfuss, Melinda Dillon, Teri Garr, Bob Balaban, Cary Guffey, and François Truffaut. It tells the story of Roy Neary, an everyday blue-collar worker in Indiana, whose life changes after an encounter with an unidentified flying object (UFO). [...] In late 1973, he developed a deal with Columbia Pictures for a science fiction film. Though Spielberg received sole credit for the script. [cont.] -1.58
Eight Oscar nominations. Close Encounters of the Third Kind. in 2017, in tribute to its 40th anniversary, the movie was given a 4K restoration of the original camera negative. Following its theatrical re-release of the director’s cut, the movie was released in 4K and Blu-ray with all three versions given the same 4K treatment. The film was nominated for 8 Oscars at the 50th Academy Awards, including Best Director, Supporting Actress (Melinda Dillon), Visual Effects, Art Direction (Joe Alves, Daniel A. Lomino, Phil Abramson), Original Music Score, Film Editing, and Sound (Robert Knudson, Robert Glass, Don MacDougall and Gene Cantamessa). The film’s only win was for Vilmos Zsigmond’s cinematography [cont.] -1.48
[’Eight nominations.’, ’Eight Oscar nominations.’, ’Eight Oscar nominations.’], [’Eight nominations, one win.’, ’Eight nominations, one win.’], [’Eight, including Best Director.’], [’Eight nominations at the 50th Academy Awards.’, ’8 nominations at the 50th Academy Awards.’], [’8 nomin ations, 1 win for cinematography.’, ’Eight Oscars, won for cinematography.’] 1.56
@@
Table 14: False positive example from (NQ development set): Passage Utility predicts a wrong answer whereas the prediction of the QA model (Gemma2-9B) is correct.
Question: What magazine, with its iconic yellow border, was first published on Sept 22, 1888? Generated Answer: The Yellow Book. Gold Answer: National Geographic.
The New Yorker. Time (magazine). plays, musicals, television programs, and literary bestsellers similar to ‘‘The New Yorker’’’s ‘‘Current Events’’ section. ‘‘Time’’ is also known for its signature red border, first introduced in 1927. The border has only been changed five times since 1927: The issue released shortly after the September 11 attacks on the United States featured a black border to symbolize mourning. However, this edition was a special ‘‘extra’’ edition published quickly for the breaking news of the event; the next regularly scheduled issue contained the red border. Additionally, the April 28, 2008, Earth Day issue, dedicated to environmental issues, contained a green border. [cont.] 1.92
The Yellow Book. The Yellow Book. The Yellow Book The Yellow Book was a British quarterly literary periodical that was published in London from 1894 to 1897. It was published at The Bodley Head Publishing House by Elkin Mathews and John Lane, and later by John Lane alone, and edited by the American Henry Harland. The periodical was priced at 5 shillings and lent its name to the ‘‘Yellow Nineties’’, referring to the decade of its operation. It was a leading journal of the British 1890s; [cont.] 0.35
The Crisis. The Colored American Magazine. The Colored American Magazine The Colored American Magazine was the first American monthly publication that covered African-American culture. The magazine ran from May 1900 to November 1909. It was initially published out of Boston by the Colored Co-Operative Publishing Company, and from 1904, forward, by Moore Publishing and Printing Company of New York. Pauline Hopkins, its most prolific writer from the beginning, sat on the board as a shareholder, was editor from 1902 to 1904, though her name was not on the masthead until 1903. [cont.] -1.27
[’The Yellow Book’, ’The Yellow Book’, ’The Yellow Book’, ’The Yellow Book’, ’The Yellow Book’, ’Time magazine.’, ’The Yellow Book’, ’The Yellow Book’, ’The Yellow Book’], [’Time magazine’] 0.50
@@
Table 15: False negative (from TQA development set): Passage Utility predicts a correct answer, and the answer by the QA model (Gemma2-9B is wrong.

### v2

@@ E.3 Prompts
The prompt we use for our QA models is shown in Table 9. Table 10 illustrates the prompt used for the p(true) baseline. Table 11 shows the prompt used for the LLM-based accuracy metric.
Appendix F Additional Results
F.1 Reference Retrieval Augmented QA Accuracy
Table 12 shows retrieval augmented QA performance for the five QA models on the development and test sets of our six tasks. We report accuracy based on token overlap (Acc) as computed in previous work, i.e., whether the gold answer is contained in the generated answer (Mallen et al., 2023; Asai et al., 2024; Xie et al., 2024) and accuracy using an LLM as a judge (AccLM). Note that AccLM is much higher than Acc across the board, which highlights the importance of using a better accuracy metric, especially when the target QA models are not fine-tuned.
F.2 Detailed Uncertainty Estimation Results
@@
Table 13 shows AURAC scores on the test sets. Table 14 shows the performance of uncertainty quantification approaches on the development set. We report AUROC and AURAC.
@@
Development Acc AccLM Acc AccLM Acc AccLM Acc AccLM Acc AccLM Acc AccLM
Gemma2-9B 0.48 0.66 0.74 0.80 0.46 0.66 0.38 0.60 0.51 0.52
Gemma2-27B 0.48 0.66 0.75 0.81 0.49 0.67 0.38 0.60 0.52 0.52
@@
NQ TQA WebQ SQuAD PopQA RefuNQ
Test Acc AccLM Acc AccLM Acc AccLM Acc AccLM Acc AccLM Acc AccLM
Gemma2-9B 0.49 0.65 0.74 0.80 0.40 0.66 0.43 0.60 0.50 0.52 0.26 0.40
Gemma2-27B 0.48 0.65 0.76 0.81 0.41 0.66 0.42 0.61 0.51 0.53 0.26 0.39
@@
Table 12: QA model performance (with $|R|=5$) on the development and test sets. We report token- and model-based accuracy (Acc and AccLM). AccLM is computed by Qwen2-72B-Instruct.
Gemma2-9B Gemma2-27B
NQ TQA WebQ SQuAD PopQA RefuNQ NQ TQA WebQ SQuAD PopQA RefuNQ
PPL 0.69 0.84 0.63 0.57 0.56 0.45 0.67 0.78 0.63 0.62 0.56 0.45
p(true) 0.75 0.85 0.71 0.63 0.71 0.53 0.76 0.89 0.73 0.67 0.70 0.54
Regular Entropy 0.70 0.84 0.63 0.59 0.57 0.46 0.69 0.79 0.64 0.61 0.58 0.45
Semantic Entropy 0.71 0.85 0.64 0.65 0.64 0.51 0.69 0.81 0.67 0.63 0.61 0.50
Utility Ranker 0.76 0.90 0.72 0.74 0.73 0.64 0.72 0.88 0.73 0.74 0.73 0.64
Llama-3.1-8B Mistral-7B-v0.3
NQ TQA WebQ SQuAD PopQA RefuNQ NQ TQA WebQ SQuAD PopQA RefuNQ
PPL 0.73 0.87 0.71 0.68 0.69 0.54 0.67 0.83 0.66 0.66 0.61 0.54
p(true) 0.76 0.89 0.75 0.70 0.71 0.59 0.71 0.86 0.70 0.69 0.65 0.56
Regular Entropy 0.73 0.87 0.72 0.70 0.69 0.56 0.67 0.84 0.69 0.65 0.62 0.51
Semantic Entropy 0.71 0.87 0.71 0.70 0.67 0.54 0.68 0.85 0.71 0.69 0.66 0.51
Utility Ranker 0.74 0.87 0.73 0.73 0.71 0.65 0.72 0.87 0.71 0.75 0.71 0.66
Table 13: AURAC values for QA models based on Gemma2-9B/27B, Llama-3.1-8B, and Mistral-7B-v0.3 on Natural Questions (NQ), TriviaQA (TQA), WebQuestions (WebQ), SQuAD, PopQA, and RefuNQ test sets.
@@
Gemma2-9B NQ TQA WebQ SQuAD PopQA NQ TQA WebQ SQuAD PopQA
@@
AvgAnsLen 0.61 0.64 0.65 0.63 0.68 0.68 0.83 0.71 0.65 0.61
Retriever Score 0.53 0.62 0.53 0.67 0.64 0.65 0.82 0.63 0.67 0.61
Utility Ranker 0.72 0.84 0.75 0.85 0.85 0.75 0.89 0.77 0.77 0.71
AUROC AURAC
Gemma2-27B NQ TQA WebQ SQuAD PopQA NQ TQA WebQ SQuAD PopQA
PPL 0.61 0.56 0.55 0.63 0.53 0.68 0.79 0.65 0.67 0.52
MSP 0.64 0.66 0.59 0.67 0.60 0.70 0.82 0.67 0.69 0.56
PMI 0.51 0.52 0.56 0.54 0.56 0.64 0.78 0.67 0.62 0.56
p(true) 0.76 0.73 0.69 0.69 0.79 0.78 0.84 0.74 0.71 0.70
Regular Entropy 0.65 0.53 0.56 0.64 0.53 0.71 0.78 0.66 0.67 0.52
Cluster Assignment 0.66 0.67 0.59 0.66 0.66 0.71 0.82 0.67 0.68 0.60
Semantic Entropy 0.64 0.67 0.59 0.68 0.66 0.69 0.82 0.68 0.69 0.60
AvgAnsLen 0.63 0.68 0.65 0.60 0.69 0.69 0.83 0.72 0.66 0.61
Retriever Score 0.56 0.60 0.51 0.69 0.65 0.67 0.81 0.64 0.71 0.62
Utility Ranker 0.73 0.75 0.72 0.84 0.87 0.75 0.86 0.75 0.78 0.73
AUROC AURAC
Llama-3.1-8B NQ TQA WebQ SQuAD PopQA NQ TQA WebQ SQuAD PopQA
@@
Semantic Entropy 0.74 0.83 0.70 0.74 0.81 0.75 0.86 0.72 0.71 0.68
AvgAnsLen 0.73 0.73 0.69 0.69 0.84 0.73 0.82 0.71 0.67 0.69
Retriever Score 0.58 0.63 0.54 0.68 0.66 0.65 0.79 0.62 0.66 0.60
Utility Ranker 0.78 0.85 0.74 0.82 0.86 0.75 0.87 0.75 0.75 0.71
AUROC AURAC
Mistral-7B-v0.3 NQ TQA WebQ SQuAD PopQA NQ TQA WebQ SQuAD PopQA
@@
AvgAnsLen 0.66 0.75 0.65 0.68 0.81 0.69 0.85 0.73 0.67 0.70
Retriever Score 0.55 0.63 0.55 0.65 0.68 0.63 0.81 0.67 0.64 0.63
Utility Ranker 0.76 0.81 0.75 0.85 0.85 0.74 0.87 0.78 0.75 0.71
Table 14: Answer uncertainty estimation for QA models Gemma2-9/27B, Llama-3.1-8B, and Mistral-7B-v0.3 on NQ, TQA, WebQ, SQuAD, and PopQA development sets. We report AUROC and AURAC.
Appendix G Examples of False Positives and Negatives
Tables 1518 illustrate the working of Passage Utility for answer uncertainty estimation. As we report AUROC scores, we do not set any correct/incorrect decision threshold; for the purpose of this discussion, we assume a decision point at 0.5 and analyze clear success and failure cases. For each example, we show the question, gold, and generated answers in the top block. Then, we show three retrieved passages with their estimated Passage Utility and a final block with ten sampled answers, their grouping into clusters, and the Cluster Assignment entropy.
Table 15 shows an example for a SQuAD question and the Llama-3.1-8B QA model. In this case, the QA model correctly answers and the Passage Utility estimate is high (i.e., indicating the answer is correct). Table 16 illustrates a case where Llama-3.1-8B’s answer is incorrect and all Passage Utilities are very low (i.e., indicating the answer is incorrect). The example from NQ in Table 17 shows a case where all Passage Utilities are low but the QA model (Gemma2-9B) answers correctly. The first passage is not useful, the second does not explicitly mention the answer but still primes the QA model to answer correctly, while the third passage mentions the answer.
@@ Appendix G Examples of False Positives and Negatives
Question: Due to increased unemployment, who mainly opposed the Shah’s regime?
Generated Answer: Millions of youth and poor migrants.
Gold Answer: Millions of youth who had migrated to the cities.
Youth who migrated to cities for construction jobs. Iran. unemployment, especially among millions of youth who had migrated to the cities of Iran looking for construction jobs during the boom years of the early 1970s. By the late 1970s, many of these people opposed the Shah’s regime and began to organize and join the protests against it. The 1979 Revolution, later known as the ‘‘Islamic Revolution’’, began in January 1978 with the first major demonstrations against the Shah. [cont.] 2.41
Unemployed and the poor. Ruhollah Khomeini. unemployment, ideological disagreement over the economy, and ‘‘international pressure and isolation’’ such as US sanctions following the hostage crisis. Due to the Iran–Iraq War, poverty is said to have risen by nearly 45% during the first 6 years of Khomeini’s rule. Emigration from Iran also developed, reportedly for the first time in the country’s history. Since the revolution and war with Iraq, an estimated ‘‘two to four million entrepreneurs, professionals, technicians, and skilled craftspeople (and their capital)’’ have emigrated to other countries. [cont.] -2.51
The National Front. Mohammad Reza Pahlavi. professors issued a public statement criticising the 1953 coup, all were dismissed from their jobs, but in the first of his many acts of ‘‘magnanimity’’ towards the National Front, Mohammad Reza intervened to have them reinstated. Mohammad Reza tried very hard to co-opt the supporters of the National Front by adopting some of their rhetoric and addressing their concerns, for example declaring in several speeches his concerns about the Third World economic conditions and poverty which prevailed in Iran, a matter that had not much interested him before. [cont.] -2.42
[’Migrants to the cities.’], [’Millions of youth who had migrated to cities.’, ’Millions of youth who migrated to cities.’], [’Cultural and religious conservatives, and the urban poor who had migrated to cities for jobs.’], [’Cultural and religious conservatives, mostly urban migrants.’, ’Cultural and religious conservatives, particularly the migrants to cities.’, ’Cultural and religious conservatives, especially those rece ntly migrated to the cities.’], [’Millions of youth and poor migrants to cities.’], [’Millions of young migrants who had moved to cities in the early 1970s.’] 1.83
@@
Question: Which company was targeted by the NAACP for not having fair practices?
Generated Answer: Target Corporation.
Gold Answer: Lockheed Aircraft Corporation.
Target Corporation. Target Corporation. of Colored People has repeatedly given Target failing grades on its annual Economic Reciprocity Initiative report card, a measure of the company’s ‘‘commitment to the African-American citizenry’’. In 2003 and 2005, the NAACP has rated Target an ‘‘F’’ on this report; in 2004, Target was rated a ‘‘D-’’. In 2006, when Target was asked why it didn’t participate in the survey again, a representative explained, ‘‘Target views diversity as being inclusive of all people from all different backgrounds, not just one group.’’ In February 2006, the National Federation of the Blind (NFB) filed a class action [cont.] -2.64
None, the NAACP was involved in the Duke lacrosse case. Reactions to the Duke lacrosse case. formed an opinion on the case. North Carolina NAACP Legal Redress Chair, Al McSurely, explained that ‘‘The NAACP stands for fair play for all parties, zealous investigation and deep concern for the survivors of racist/sexist attacks.’’ At the same time, some have criticized the NAACP for making statements that portrayed the players as racist despite evidence to the contrary, using the case to promote the group’s cause, and implying guilt. McSurely stated that ‘‘[w]ithin five minutes, the men threatened the women with racial and misogynist verbal assaults, [cont.] -2.77
Philadelphia Transit Company (PTC). Philadelphia transit strike of 1944. (PRTEU), Frank Carney, proved to be equally reticent and claimed that he was not authorized by the union members to consider a request to allow promotions of black employees. The black PTC employees enlisted the help of the NAACP and started lobbying the federal authorities, particularly the Fair Employment Practices Commission (FEPC), to intervene. The Fair Employment Practices Commission, created by an executive order of the President in 1941, was charged with ensuring non-discrimination employment practices by government contractors. [cont.] -2.60
@@
Table 16: True positive example (from SQuAD development set): Passage Utility predicts the answer provided by the QA model (Llama-3.1-8B) which is incorrect.
Question: What do Alfred Newman, John Williams, Johnny Green, Andr Previn, John Barry and Alan Menken have in common?
Generated Answer: They are all Oscar-winning film composers.
Gold Answer: Winning the Academy award for Best Original Score.
They all won multiple Academy Awards for film scores. Alan Menken. his other film musicals were also nominated: Alan Menken Alan Irwin Menken (born July 22, 1 949) is an American musical theatre and film score composer and pianist. Menken is best known for his scores for films produced by Walt Disney Animation Studios. His scores for "The Little Mermaid" (1989 ), "Beauty and the Beast" (1991), "Aladdin" (1992), and "Pocahontas" (1995) have each won him two Academy Awards. He also composed the scores for "Little Shop of Horrors" (1986), "Newsies" (1992), "The Hunchback of Notre Dame" (1996), "Hercules" (1997), "Home on the Range" (2004), "Enchanted" (2007), "Tangled" (2010), and "Sausage Party" (2016) -2.22
They are all prolific Oscar winners in music categories. Alan Menken. hi s work on musical theatre works for Broadway and elsewhere. Some of these are based on his Disney films, but other stage hits include "Little Shop of Horrors" (1982), "A Christmas Carol" (1994) and "S ister Act" (2009). Menken has collaborated with such lyricists as Howard Ashman, Tim Rice, Glenn Slater, Stephen Schwartz and David Zippel. With eight Academy Award wins (four each for Best Score and Best Song), Menken is the second most prolific Oscar winner in the music categories after Alfred Newman, who has nine Oscars. He has also won eleven Grammy Awards, a Tony Award and other honors. -2.30
They are all Academy Award-winning composers. Alan Menken. the Beast", with the so ngs from the 1991 film and new material by Menken and Rice. , Menken is collaborating on writing new songs with Benj Pasek and Justin Paul for a live-action film version of "Aladdin" and is also working with Lin-Manuel Miranda on new music for a live-action film adaptation of "The Little Mermaid". Menken will also once again be working with Stephen Schwartz to write new songs for D̈isenchanted,̈ the s equel to "Enchanted". With eight Academy Awards (four each for best score and best song), only composer Alfred Newman (nine wins) and Walt Disney (22 wins) have -2.31
[’They are all Academy Award-winning composers.’, ’They are all Academy Award-winning composers.’, ’They are all Oscar-winning film composers.’, ’They are all Oscar-winning film composers.’], [’They are all multiple-time Oscar winners for music.’, ’They are prolific Academy Award-winning composers.’], [’They are all multi-Oscar winning film composers.’, ’They are all multiple Oscar-winning composers.’, ’They are prolific Oscar-winning film composers.’], [’They are all prolific film music composers.’] 1.28
@@
Table 17: False positive example (from TQA development set): Passage Utility predicts a wrong answer whereas the prediction of the QA model (Gemma2-9B) is correct.
Question: One Step Behind’ is a tribute act to which band?
Generated Answer: Furniture.
Gold Answer: North London Invaders.
Def L eppard. Two Steps Behind. Two Steps Behind "Two Steps Behind" is a 1993 song by English hard rock band Def Leppard from their album "Retro Active" and "Last Action Hero" Soundtrack. It reached #5 on the "Billboard" Hot Mainstream Rock Tracks chart, and #12 on the B̈illboardḦot 100. Though the band did manage to chart a few more songs in the following years, this is generally considered to be the band’s last major hit in the US. In the 1993 "Metal Edge" Readers’ Choice Awards, the song was voted "Song of the Year" and "Best Song From a Movie Soundtrack.". Def Leppard have 3.29
Def Leppard. Two Steps Behind. and a live footage. The video was aired on August 19 93. Filipino-Chinese singer Rachelle Ann Go covered the song for her 2007 album "Obsession". Two Steps Behind "Two Steps Behind" is a 1993 song by English hard rock band Def Leppard from their album "Retro Active" and "Last Action Hero" Soundtrack. It reached #5 on the "Billboard" Hot Mainstream Rock Tracks chart, and #12 on the "Billboard" Hot 100. Though the band did manage to chart a few more songs in the following years, this is generally considered to be the band’s last major hit in the US. In the 2.42
Split Enz. One Step Ahead (Split Enz song). unavailable to Australasian markets until 2007 when it became available on iTunes). The video clip to "One Step Ahead" has keyboardist Eddie Rayner performing "Marche sur place", the pantomime illusion walk created by Decroux and Barrault (seen in the 1945 French film Child ren of Paradise) that is the technique Michael Jackson would base his moonwalk on in 1983. One Step Ahead (Split Enz song) "One Step Ahead" is a 1980 song by New Zealand art rock group Split Enz from the ir studio album "Waiata". The song continued the group’s success in their move towards their own version of new wave -2.93
[’Furniture’, ’Furniture’, ’Furniture’, ’Furniture’, ’Furniture’, ’Furniture’, ’Furniture’, ’Furniture’, ’Furniture’, ’Furniture’] 0
@@
Table 18: False negative (from TQA development set): Passage Utility predicts a correct answer, and the answer by the QA model (Gemma2-9B) is wrong.